OpenStudy (anonymous):

Frank can type a report in 3 hrs and James takes 4hrs .how long will it take two of them typing together?

OpenStudy (dumbcow):

you have to add the 2 rates: Frank is working at rate of 1/3 report per hour James at rate of 1/4 report per hour --> 1/3 + 1/4 = 4/12 + 3/12 = 7/12 together they work at rate 7/12 report per hour time to finish report is then 12/7 hours

OpenStudy (anonymous):

Thnk you